ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

17. Control Theory and Optimization

1991 MS Classification: Wessäly, Roland, Konrad Zuse Zentrum fuer Informationstechnik Berlin New inequalities for a capacitated survivable network design problem In this talk we present a new class of inequalities for a polytope associated with a capacitated survivable network design problem. In the underlying problem, which came up in a practical application with a german mobilephone company, we wish to choose a capacity for each available link of the network such that the overall cost is minimal, the estimated communication demands can be routed and the network satisfies certain survivability requirements.